Circular No. 8841

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 K6 (McNAUGHT) R. H. McNaught reports his discovery of another comet on 0.5-m Up-

psala Schmidt telescope images obtained in the course of the Siding Spring Survey (discovery observation tabulated below), the object displaying a marginal coma of diameter ∼ 6′′. Additional images by McNaught on May 30.8 UT show an almost, but-not-quite-asteroidal appearance in excellent seeing, the coma diameter given as 4′′ (FWHM).

2007 UT  α_{2000}  δ_{2000}  Mag.
May 27.79676  22^{h}50^{m}33.^{s}10  −42^{o}40′59′′.9  18.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-K82.

T = 2007 June 21.950 TT  ω = 334.^{o}978
Ω = 298.186 }2000.0
q = 3.46556 AU  i = 105.095

COMET C/2007 K5 (LOVEJOY) Additional precise astrometry of this comet (cf. IAUC 8840) has been

published on MPEC 2007-K80, where the following preliminary parabolic orbital elements also appear:

T = 2007 Apr. 26.977 TT  ω = 250.^{o}115
Ω = 191.798 }2000.0
q = 1.12113 AU  i = 65.141

COMET C/2007 JA_{21} (LINEAR) Further to IAUC 8837, J. Young reports that his CCD images taken

on May 14.29–14.36 UT with the Table Mountain 0.61-m f/16 Cassegrain reflector show a coma of diameter 5′′–6′′ that was slightly elongated along the comet’s path of motion, with a faint fan-shaped tail extending 20′′ spanning p.a. 315^{o}–25^{o}. Improved orbital elements from MPEC 2007-K78:

T = 2006 Nov. 10.980 TT  ω = 93.^{o}308
Ω = 65.556 }2000.0
q = 5.35972 AU  i = 89.825
2007 May 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green